OpenStudy (anonymous):

Find the exact value of cos(u+v) given that sin u =7/25 and cos v = -12/13 (both u and v are in Quadrant 2). I am stuck with getting 288/325.

OpenStudy (anonymous):

-209/325

OpenStudy (anonymous):

7/25-12/13=-209/325

OpenStudy (anonymous):

should be 253/325

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

My answer key said that was the answer. You could be correct. :) My book uses Cos (v-u).

OpenStudy (anonymous):

\[\frac{ 7 }{ 25 }*\frac{ 13 }{ 13 }-\frac{ 12 }{ 13 }*\frac{ 25 }{ 25 }\]

OpenStudy (anonymous):

cos(-209/325)=0.8002

OpenStudy (anonymous):

my book does it differently...but, I see what you did. You are likely right.

OpenStudy (anonymous):

what grade is this?

OpenStudy (anonymous):

Trig

OpenStudy (anonymous):

(-24/25)(-12/13)(7/25)(18/13)

OpenStudy (anonymous):

?/325

OpenStudy (anonymous):

where did all of these numbers come from